# A decade of the fast-varying ionospheric and magnetospheric magnetic fields

### v1

@@ 4 Model construction and selection
The magnetic field model is parametrised in three sets of source-specific Gauss coefficients: $g_{nm}^{\mathrm{int}}(t)$ and $h_{nm}^{\mathrm{int}}(t)$ for the field induced within the Earth, $q_{nm}^{\mathrm{ion}}(t)$ and $s_{nm}^{\mathrm{ion}}(t)$ for the ionospheric field, and $q_{nm}^{\mathrm{mag}}(t)$ and $s_{nm}^{\mathrm{mag}}(t)$ for the magnetospheric field (Eq. 13-14). Denoting these Gauss coefficients in the $i$-th time bin jointly as $\mathbf{m}_{i}$, and the observatory and satellite data jointly as $\mathbf{d}_{i}$, the model estimation procedure in each time bin can be formulated as solving the following optimisation problem
@@
where $\mathbf{W}_{i}$ is the data weight matrix for the data in the $i$-th time bin, $\mathbf{G}_{i}$ is the Gauss matrix, and $L(\cdot)$ is a generic misfit function. We assume independent sampling of each data point and assign uniform variance to each observation, hence $\mathbf{W}=\mathbf{I}$. The Gauss matrix $\mathbf{G}_{i}$ is the design matrix for the regression problem, and can be obtained by evaluating the gradients of the potentials in Eq. (13-14) at observation locations.
@@
Since no temporal behaviour is imposed on any of the SH modes, all the modes and their coefficients with a spherical degree lower than the truncation degree are treated equally in Eq. (13-14). Since spherical harmonics form an irreducible representation of the 3-D rotation group, there is no difference between using different coordinate systems in our formulation. We chose to work in geocentric geographic coordinates to simplify integration with the existing dataset, as well as with other ground-based observations that are naturally provided in this coordinate system.
@@
The truncation degree for the spherical harmonics is selected based on two arguments. First, the local time coverage of the satellites gives a theoretical upper bound on the maximum azimuthal wavenumber. Our dataset contains data from six satellites, two of which are at the same local time (Swarm A and C). With an approximate orbital period of 90 minutes, the satellite orbits within a $3$-hr time bin cover a maximum of 10 local times. Analogous to the Nyquist sampling theorem, this indicates that the highest azimuthal wavenumber field that can be sampled is $m=5$. Beyond this azimuthal wavenumber, even if the combined external field can be estimated from the observatory dataset, it can no longer be convincingly separated into ionospheric and magnetospheric contributions without aliasing by the satellites. As we are using $N=M$, this limits the truncation SH degree to $N\leq 5$.
@@
Given $N\leq 5$, the specific parametrisation is further selected using $K$-fold cross-validation (CV), a technique that provides an estimate on the generalisation error of the model. In a $K$-fold CV, the full dataset to be used for parameter estimation is split into $K$ parts. Model parameters are estimated in each fold using $K-1$ parts of the datasets, while the remaining part serves as the validation set for testing how the estimated model performs on unseen data. The average of the $K$ validation errors then yields the proxy for the generalisation error (Arlot & Celisse, 2010). While a pure extension of the model parametrisation will always reduce the total data misfit (provided global minimum is reached), the generalisation error will stagnate or deteriorate when the model is being increasingly fitted to the noise rather than the true signal, and hence provides means to detect overfitting.
@@
We performed a 5-fold CV on the combined dataset in each time bin, for SH truncation degrees ranging from $3$ to $7$. The coefficient of determination $R^{2}$ is then computed for the ground observatory and satellite datasets separately, as performance scores for each of the model parametrisations. Statistics of the CV $R^{2}$ score shows that while the estimated model seems to perform better as the truncation degree increases on the satellite dataset (blue line in Fig. 5), the CV score for the ground observatory dataset stagnates at $N=5$ and then decreases for $N=7$. This observation is consistent with the theoretical argument that $N>5$ is beyond the spatial resolution of our dataset (blue line in Fig. 4). A more careful investigation shows that a model with induced and magnetospheric fields parametrised to $N=4$ but ionospheric field extended to $N=5$ already explains most of the variance of the data and is located on the CV score plateau of $N=5$. Compared to this model, the 25% more free parameters in the $N=5$ model only yields a marginal increase of 0.002 in the CV score. This implies that the extra degrees of freedom in the $N=5$ model does not yield a significant gain in explaining the data, while increasing the risk of overfitting. Based on these observations, we adopted the parametrisation in which the induced and magnetospheric fields are expanded to $N=4$, and the ionospheric field is expanded to $N=5$. For each time bin, Gauss coefficients $\mathbf{m}_{i}$ are estimated by solving Eq. (15) using a robust Huber loss as the loss function $L(\cdot)$.
@@
Figure 4: The coefficient of determination $R^{2}$ from fitting the whole ground observatory dataset (red) and from CV (blue) for models estimated using a SH truncation degrees from $3$ to $6$. The square markers show the median of the $R^{2}$ scores, while the upper and lower error bar limits show the 1st and 3rd quantile, respectively.
@@ 5.1 The role of ionosphere in the model
By estimating coefficients in each time bin, we obtained the 10-year time series of Gauss coefficients describing the magnetospheric field, the ionospheric field and the mantle induced counterpart. The introduction of the ionospheric field in our approach, as described in Section 2, is the major reason why all three fields can be characterised individually. While inclusion of the ionosphere is dictated by reality, it results in additional model parameters and requires the sheet current approximation. Hence, a statistical experiment is needed to justify that adding the ionosphere indeed results in a better model.
To this end, we constructed another model with the same model parametrisation, but without the ionosphere layer. This model effectively attributes all variations in the satellite and observatory data to a common internal source (the mantle induced currents) and a common external source (the magnetospheric currents). Comparing the results from the reduced model (model without the ionosphere) to the results we obtained using the complete model with the ionosphere provides a data-based criterion for justifying the addition of the ionospheric component into our model.
Naturally, introducing an ionospheric field is able to explain more variance in the data simply by expanding the parameter space. The data fit alone is hence not a good indicator of the quality of a model, as the model can also achieve lower misfit by fitting the noise in the dataset. To objectively compare the quality of the model and rule out the possibility of overfitting, we once again employ the CV approach. Using a 5-fold CV, we see that there is a drastic improvement in the CV averaged $R^{2}$ score (Fig. 6 lower panel) when an ionosphere layer is included. The parametrisation with an ionosphere is thus not only better at explaining the data variance overall, but can explain the unseen observations better. This is a strong indication that the model parametrisation with an ionosphere provides better descriptions of the geomagnetic field.
@@
As is evident from different $R^{2}$ values, introducing the ionosphere in our model parametrisation certainly changes the data fit. It is however unclear how the introduction of the ionospheric field affects the existing field estimates - that is, how the estimates on the induced and magnetospheric magnetic fields are affected by the introduction of the ionosphere. To further quantify the impact of the ionospheric layer on the magnetospheric field coefficient estimates, we turn to two statistical overviews of our field estimates. First, as an overview of the difference in the Gauss coefficients, we compute the root-mean-square (RMS) differences between the magnetospheric field coefficient time series estimated with and without the ionosphere, defined as
$\mathrm{RMSD}_{nm}^{\mathrm{mag}}=\left\{\begin{aligned} \sqrt{\frac{1}{K}\sum_{k=1}^{K}\left[q_{nm}^{\mathrm{mag}}(t_{k})-\widetilde{q}_{nm}^{\mathrm{mag}}(t_{k})\right]^{2}},\quad m\geq 0\\ \sqrt{\frac{1}{K}\sum_{k=1}^{K}\left[s_{n|m|}^{\mathrm{mag}}(t_{k})-\widetilde{s}_{n|m|}^{\mathrm{mag}}(t_{k})\right]^{2}},\quad m<0.\end{aligned}\right.$ (16)
The absolute RMS differences, together with their relative values using the full model estimates as a reference, are shown in Fig. 7 for different spherical harmonics. The notation $\widetilde{f}$ here denotes the field coefficients $f$ estimated in the reduced model, without an ionosphere. In the notation for the RMS differences, a negative $m$ is used to denote sine modes, whereas the non-negative $m$ is reserved for cosine modes. For $q_{10}^{\mathrm{mag}}$, the RMS difference with and without an ionosphere is $2.4$nT during storm time ($Kp\geq 5$), and as low as $0.86$ during quiet time ($Kp\leq 2$). Due to the strong magnetospheric $P_{1}^{0}$ signal, these differences translate to merely $3.8\%$ and $4.7\%$ during storm time and quiet time, respectively. The coherence of the two model estimates also extends to other spherical harmonics, such as the $q_{21}^{\mathrm{mag}}$ component. For this mode, the estimates with and without an ionosphere have an RMS difference of $2.1$nT during storm time and $1.0$nT during quiet time. Although the RMS differences are always smaller during quiet times, we observe that the relative differences are typically higher (Fig. 7), due to the overall weaker magnetospheric field and greater contributions from the ionospheric field.
@@
Figure 9: Relative RMS difference between the SH coefficient time series estimated using the models with and without the ionosphere. The rows and columns are the same as in Fig. 7
@@ 5.2 Time series of external and induced coefficients
In addition to the magnetospheric field, our model also simultaneously estimates the ionospheric field Gauss coefficients. The $q_{21}^{\mathrm{ion}}$ spatial mode (Fig. 13) exhibits a large amplitude in the diurnal band in the geographic frame, reflecting a standing mode in the solar-fixed frame, an important constituent of the Sq variation (Malin, 1973; Schmucker, 1999). The diurnal variation of $q_{21}^{\mathrm{ion}}$ is very clear in the reconstructed time series not only for magnetic quiet times (Fig. 13 lower panel), but around the geomagnetic storms as well (Fig. 13 upper panel). This suggests that the Sq variation is a robust feature that is modulated by the magnetosphere even during geomagnetic storms, e.g. during the recovery phases, although its amplitude is dwarfed by a much larger $q_{10}^{\mathrm{mag}}$ magnetospheric signal (Fig. 11). However, we observe a considerable deviation from the quiet time daily variation during the storm’s main phase on the 8 September (Fig. 13 upper panel). Interestingly, this deviation is anti-correlated with the magnetospheric mode $q_{21}^{\mathrm{mag}}$ (Fig. 12). This anti-correlation may represent interaction between the ionosphere and magnetosphere current systems during the main phase of a geomagnetic storm.
@@
As has been observed in Figs. 9-10, the induced field estimate is more susceptible to the presence of the ionospheric layer to the model. This is confirmed in the time series in Figs. 14 and 15. The discrepancy is especially pronounced in spatial modes where the ionosphere is comparatively energetic, e.g. the $P_{2}^{1}$ SH mode. During the magnetic quiet time (cf. Fig. 15, lower panel), the amplitude of the induced field estimate in the reduced model is less than 50% of that estimated using the full model with an ionosphere. We interpret this systematically smaller amplitude as an imprint of the missing ionospheric field on the induced field. When the ionosphere is neglected, the internal contribution to the satellite data which should have been attributed to the ionosphere is forced into the induced mantle coefficients. This, in turn, may weaken the induction signal when the induced field is correlated to the ionospheric signal (Eq. 14). The difference in the model estimates with and without the ionosphere is not only in the amplitude, but also in the phase. The phase change can already be observed in $g_{21}^{\mathrm{int}}$, but it is even more pronounced in higher SH degrees. The semi-diurnal oscillation of $h_{32}^{\mathrm{int}}$, for instance, appears to be $\sim 3$ hours ahead in the model without an ionosphere compared to the full model estimate (Fig. 16).
@@
Figure 14: Time series of the mantle induced Gauss coefficient $g_{10}^{\mathrm{int}}(t)$ around the 2017 September geomagnetic storm (upper panel) and during magnetic quiet time, in January / February 2018 (lower panel). The plots show our full model estimates (brown lines) and the estimates with no ionosphere parametrisation (orange).
@@
Figure 15: Time series of the mantle induced Gauss coefficient $g_{21}^{\mathrm{int}}(t)$ around the 2017 September geomagnetic storm (upper panel) and during magnetic quiet time, in January / February 2018 (lower panel).
@@
Figure 16: Time series of the mantle induced Gauss coefficient $h_{32}^{\mathrm{int}}(t)$ during magnetic quiet time, in February 2018.
5.3 Fourier spectrum of external fields

### v2

@@ 4 Model construction and selection
The magnetic field model is parametrised in three sets of source-specific Gauss coefficients: $g_{nm}^{\mathrm{int}}(t)$ and $h_{nm}^{\mathrm{int}}(t)$ for the internally induced field, $q_{nm}^{\mathrm{ion}}(t)$ and $s_{nm}^{\mathrm{ion}}(t)$ for the ionospheric field, and $q_{nm}^{\mathrm{mag}}(t)$ and $s_{nm}^{\mathrm{mag}}(t)$ for the magnetospheric field (Eq. 13-14). Denoting these Gauss coefficients in the $i$-th time bin jointly as $\mathbf{m}_{i}$, and the observatory and satellite data jointly as $\mathbf{d}_{i}$, the model estimation procedure in each time bin can be formulated as solving the following optimisation problem
@@
where $\mathbf{W}_{i}$ is the data weight matrix for the data in the $i$-th time bin, $\mathbf{G}_{i}$ is the Gauss matrix, and $L(\cdot)$ is a generic misfit function. We assume independent sampling of each data point and assign uniform variance to each observation, hence $\mathbf{W}=\mathbf{I}$. The Gauss matrix $\mathbf{G}_{i}$ is the design matrix for the regression problem, and can be obtained by evaluating the gradients of the potentials in Eqs. (13-14) at observation locations.
@@
Since no temporal behaviour is imposed on any of the SH modes, all the modes and their coefficients with a spherical degree no greater than the truncation degree are treated equally in Eqs. (13-14). Since spherical harmonics form an irreducible representation of the 3-D rotation group, there is no difference between using different coordinate systems in our formulation. We chose to work in geocentric geographic coordinates to simplify integration with the existing dataset, as well as with other ground-based observations that are naturally provided in this coordinate system.
@@
The truncation degree for the spherical harmonics is selected based on two arguments. First, the local time coverage of the satellites gives a theoretical upper bound on the maximum azimuthal wavenumber. Our dataset contains data from six satellites in polar orbits, which, with an approximate orbital period of 90-100 minutes, cover 24 local times within a $3$-hr time bin (each satellite covers two local times per orbit, and completes two orbits per time bin). However, among these satellites, two pairs always share the same local times as they are very close in orbits (Swarm A and C, Grace-FO 1 and 2), and Swarm B satellite also evolves to the same orbital plane as Swarm A and C around 2021. Therefore, the satellites at best sample the magnetic field at 12 local times within one time bin. According to the Nyquist sampling theorem, the field that can be reconstructed without aliasing is limited to azimuthal wavenumber $m<6$, even with uniform sampling. Beyond this azimuthal wavenumber, even if the combined external field can be estimated from the observatory dataset, it can no longer be convincingly separated into ionospheric and magnetospheric contributions without being aliased by the satellites. As we are using $N=M$, this limits the truncation SH degree to $N<6$.
@@
Given $N<6$, the specific parametrisation is further selected using $K$-fold cross-validation (CV), a technique that provides an estimate on the generalisation error of the model. In a $K$-fold CV, the full dataset to be used for parameter estimation is split into $K$ parts. Model parameters are estimated in each fold using $K-1$ parts of the datasets, while the remaining part serves as the validation set for testing how the estimated model performs on unseen data. The average of the $K$ validation errors then yields the proxy for the generalisation error (Arlot & Celisse, 2010). While a pure extension of the model parametrisation will always reduce the total data misfit (provided global minimum is reached), the generalisation error will stagnate or deteriorate when the model is being increasingly fitted to the noise rather than the true signal, and hence provides means to detect overfitting.
We performed a 5-fold CV on the combined dataset in each time bin, for SH truncation degrees ranging from $3$ to $7$. The performance of the model is evaluated using the coefficient of determination, defined as
$R^{2}=1-\frac{\|\mathbf{d}-\hat{\mathbf{d}}\|_{2}^{2}}{\|\mathbf{d}\|_{2}^{2}-N_{d}\overline{d}^{2}}$ (16)
@@
where $\mathbf{d}$ is the observed data vector, $\hat{\mathbf{d}}$ is data predicted by a model, $\overline{d}$ is the arithmetic mean of $\mathbf{d}$, and $N_{d}=\dim\mathbf{d}$ the dimension of the data. For each of the model parametrisations, the coefficients of determination $R^{2}$ are computed on the validation set of the ground observatory and satellite datasets separately and averaged over the CV folds. Statistics of the CV $R^{2}$ score shows that while the estimated model seems to perform better as the truncation degree increases on the satellite dataset (blue line in Fig. 5), the CV score for the ground observatory dataset stagnates at $N=5$ and then decreases for $N=7$ (blue line in Fig. 4). This observation is consistent with the theoretical argument that SH modes with degrees $>5$ cannot be resolved with the current data. A more careful investigation shows that a model with induced and magnetospheric fields parametrised to $N=4$, but ionospheric field extended to $N=5$ already explains most of the variance in the data and is located on the CV score plateau of $N=5$. Compared to this model, the 25% more free parameters in the $N=5$ model only yields a marginal increase of 0.002 in the CV score. This implies that the extra degrees of freedom in the $N=5$ model does not yield a significant gain in explaining the data, while increasing the risk of overfitting. Based on these observations, we adopted the parametrisation in which the induced and magnetospheric fields are expanded to $N=4$, and the ionospheric field is expanded to $N=5$. For each time bin, Gauss coefficients $\mathbf{m}_{i}$ are estimated by solving Eq. (15) using a robust Huber loss as the loss function $L(\cdot)$.
@@
Figure 4: The coefficient of determination $R^{2}$ from fitting the whole ground observatory dataset (red) and from CV (blue) for models estimated using SH truncation degrees from $3$ to $7$. The square markers show the median of the $R^{2}$ scores, while the upper and lower error bar limits show the 1st and 3rd quantiles, respectively.
@@ 5.1 The role of ionosphere in the model
By estimating coefficients in each time bin, we obtained the 10-year time series of Gauss coefficients describing the magnetospheric field, the ionospheric field, and the induced counterpart. The introduction of the ionospheric field in our approach, as described in Section 2, is the major reason all three fields can be characterised individually. While the inclusion of the ionosphere is dictated by reality, it results in additional model parameters and requires the sheet current approximation. Hence, a statistical experiment is needed to justify that adding the ionosphere indeed results in a better model.
To this end, we constructed another model with the same model parametrisation, but without the ionosphere layer. This model effectively attributes all satellite and observatory data variations to a common internal source (internally induced currents) and a common external source (magnetospheric currents). Comparing the results from the reduced model (model without the ionosphere) to the results we obtained using the complete model with the ionosphere provides a data-based criterion for justifying the addition of the ionospheric component into our model.
Naturally, introducing an ionospheric field is able to explain more variance in the data simply by expanding the parameter space. The data fit alone is, hence, not a good indicator of the quality of a model, as the model can also achieve lower misfits by fitting noise. To objectively compare the quality of the model and rule out the possibility of overfitting, we once again employ the CV approach. Using a 5-fold CV, we see that there is a drastic improvement in the CV averaged $R^{2}$ score (Fig. 6) when an ionosphere layer is included. The parametrisation with an ionosphere is thus not only better at explaining the data variance overall, but can explain the unseen observations better. This strongly indicates that the model parametrisation with an ionosphere provides better descriptions of the geomagnetic field.
@@
As is evident from different $R^{2}$ values, introducing the ionosphere in our model parametrisation certainly changes the data fit. It is however unclear how the introduction of the ionospheric field affects the existing field estimates - that is, how the estimates of the induced and magnetospheric magnetic fields are affected by the introduction of the ionosphere. To further quantify the impact of the ionospheric layer on the magnetospheric field coefficient estimates, we turn to two statistical overviews of our field estimates. First, as an overview of the difference in the Gauss coefficients, we compute the root-mean-square (RMS) differences between the magnetospheric field coefficient time series estimated with and without the ionosphere, defined as
$\mathrm{RMSD}_{nm}^{\mathrm{mag}}=\left\{\begin{aligned} \sqrt{\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}\left[q_{nm}^{\mathrm{mag}}(t_{i})-\widetilde{q}_{nm}^{\mathrm{mag}}(t_{i})\right]^{2}},\quad m\geq 0\\ \sqrt{\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}\left[s_{n|m|}^{\mathrm{mag}}(t_{i})-\widetilde{s}_{n|m|}^{\mathrm{mag}}(t_{i})\right]^{2}},\quad m<0\end{aligned}\right.$ (17)
where $N_{t}$ is the total number of time bins. The absolute RMS differences, together with their relative values using the full model estimates as a reference, are shown in Fig. 7 for different spherical harmonics. The notation $\widetilde{f}$ here denotes the field coefficients $f$ estimated in the reduced model, without an ionosphere. In the notation for the RMS differences, a negative $m$ is used to denote sine modes, whereas the non-negative $m$ is reserved for cosine modes. For $q_{10}^{\mathrm{mag}}$, the RMS difference with and without an ionosphere is $2.4$nT during storm time ($Kp\geq 5$), and as low as $0.86$nT during quiet time ($Kp\leq 2$). Due to the strong magnetospheric $P_{1}^{0}$ signal, these differences translate to merely $3.8\%$ and $4.7\%$ during storm time and quiet time, respectively. The similarity of the two model estimates also extends to other spherical harmonics, such as the $q_{21}^{\mathrm{mag}}$ component. For this mode, the estimates with and without an ionosphere have an RMS difference of $2.1$nT during storm time and $1.0$nT during quiet time. Although the RMS differences are always smaller during quiet times, we observe that the relative differences are typically higher (Fig. 7), due to the overall weaker magnetospheric field and greater contributions from the ionospheric field.
@@
Figure 9: RMS difference for Gauss coefficients of the internally induced field estimated using the models with and without the ionosphere. The rows and columns are the same as in Fig. 7.
@@ 5.2 Time series of external and induced coefficients
In addition to the magnetospheric field, our model also simultaneously estimates the ionospheric field Gauss coefficients. The $q_{21}^{\mathrm{ion}}$ spatial mode (Fig. 13) exhibits a large amplitude in the diurnal band in the geographic frame, reflecting a standing mode in the solar-fixed frame, an important constituent of the Sq variation (Malin, 1973; Schmucker, 1999). The diurnal variation of $q_{21}^{\mathrm{ion}}$ is very clear in the reconstructed time series not only for magnetic quiet times (Fig. 13 lower panel), but around the geomagnetic storms as well (Fig. 13 upper panel). This suggests that the Sq variation is a robust feature that is modulated by the magnetosphere. For instance, Sq dominant structure quickly recovers after the storm’s peak, although its amplitude is dwarfed by a much larger $q_{10}^{\mathrm{mag}}$ magnetospheric signal (Fig. 11). However, we observe a considerable deviation from the quiet time daily variation during the storm’s main phase (Fig. 13 upper panel). Interestingly, this deviation is anti-correlated with the magnetospheric mode $q_{21}^{\mathrm{mag}}$ (Fig. 12). This anti-correlation may represent an interaction between the ionosphere and magnetosphere current systems during the main phase of a geomagnetic storm.
@@
As has been observed in Figs. 9-10, the induced field estimate is more susceptible to the presence of the ionospheric layer to the model. This is confirmed in the time series in Figs. 14 and 15. The discrepancy is especially pronounced in spatial modes where the ionosphere is comparatively energetic, e.g. the $P_{2}^{1}$ SH mode. During the magnetic quiet time (cf. Fig. 15, lower panel), the amplitude of the induced field estimate in the reduced model is less than 50% of that estimated using the full model with an ionosphere. We interpret this systematically smaller amplitude as an imprint of the missing ionospheric field on the induced field. When the ionosphere is neglected, the internal contribution to the satellite data which should have been attributed to the ionosphere is forced into the induced coefficients. This, in turn, may weaken the induction signal when the induced field is correlated to the ionospheric signal (Eq. 14). The difference in the model estimates with and without the ionosphere is not only in the amplitude, but also in the phase. The phase change can already be observed in $g_{21}^{\mathrm{int}}$, but it is even more pronounced in higher SH degrees. The semi-diurnal oscillation of $h_{32}^{\mathrm{int}}$, for instance, appears to be $\sim 3$ hours ahead in the model without an ionosphere compared to the full model estimate (Fig. 16).
@@
Figure 14: Time series of the internally induced Gauss coefficient $g_{10}^{\mathrm{int}}(t)$ around the 2017 September geomagnetic storm (upper panel) and during magnetic quiet time, in January / February 2018 (lower panel). The plots show our full model estimates (brown lines) and the estimates with no ionosphere parametrisation (orange).
@@
Figure 15: Time series of the internally induced Gauss coefficient $g_{21}^{\mathrm{int}}(t)$ around the 2017 September geomagnetic storm (upper panel) and during magnetic quiet time, in January / February 2018 (lower panel).
@@
Figure 16: Time series of the internally induced Gauss coefficient $h_{32}^{\mathrm{int}}(t)$ during magnetic quiet time, in February 2018.
5.3 Fourier spectrum of external and induced fields